4. Some equations that are not separable can be made separable by an appropriate substitution. Differential equations of theplease show all steps
Answer #1

given that 44 +4 tys Differential equation Solve the above Solution - from o dt ty Aboue Differential ean is homogeneous firsTake ontilag we have (64+1jY4 = ct 64+1 = c4+4 - sinu ylt) = twit) . Then from @qn 3 probablement (glt)) #1 = citt 441t) +4 =